OpenStudy (clara1223):

find the limit as x approaches 1 of abs(x-1)/x-1

OpenStudy (clara1223):

\[\lim_{x \rightarrow 1}\frac{ \left| x-1 \right| }{ x-1 }\]

random231 (random231):

okay what is the first thing you check for when you solve a limit?

OpenStudy (dinamix):

its 1 or -1

OpenStudy (clara1223):

@random231 you plug it in to see if you get 0/0

random231 (random231):

nope dinamix i dont think you got it!

OpenStudy (clara1223):

@random231 and if so you have to simplify

random231 (random231):

no clara the first thing you have to check whether the limit exists or not!

OpenStudy (jdoe0001):

\(\large { \lim\limits_{x\to 1}\ \cfrac{|x-1|}{x-1}\implies \begin{cases} \lim\limits_{x\to 1^{\color{red}{ +}}}\ \cfrac{+(x-1)}{x-1} \\ \quad \\ \lim\limits_{x\to 1^{\color{red}{ -}}}\ \cfrac{-(x-1)}{x-1} \end{cases} }\)

OpenStudy (dinamix):

\[\lim_{x \rightarrow 1^+} = \frac{ x-1 }{ x-1 } = 1 \]

random231 (random231):

that is you put in a value immediate to the left of 1 in one case and to the right in the other case. if both of them give the same value then the limit exists.

OpenStudy (dinamix):

yup @ what i said @random231 its 1 and -1

OpenStudy (dinamix):

the answer

random231 (random231):

nope there cant be two values

OpenStudy (clara1223):

@dinamix if thats the case then doesn't the limit not exist? if theyre two different numbers?

random231 (random231):

exactly

OpenStudy (jdoe0001):

notice the values found for the two one-sided limits they differ thus the two-sided limit of \(\lim\limits_{x\to 1}\ \cfrac{|x-1|}{x-1}\impliedby \textit{does not exist}\)

OpenStudy (dinamix):

 this is graph of function now we understand everything i think i am right about my graph draw it @clara1223 , @random231